### v1

# OmniJet-: The first cross-task foundation model for particle physics

Abstract
Foundation models are multi-dataset and multi-task machine learning methods that once pre-trained can be fine-tuned for a large variety of downstream applications. The successful development of such general-purpose models for physics data would be a major breakthrough as they could improve the achievable physics performance while at the same time drastically reduce the required amount of training time and data.
We report significant progress on this challenge on several fronts. First, a comprehensive set of evaluation methods is introduced to judge the quality of an encoding from physics data into a representation suitable for the autoregressive generation of particle jets with transformer architectures (the common backbone of foundation models). These measures motivate the choice of a higher-fidelity tokenization compared to previous works. Finally, we demonstrate transfer learning between an unsupervised problem (jet generation) and a classic supervised task (jet tagging) with our new OmniJet- model. This is the first successful transfer between two different and actively studied classes of tasks and constitutes a major step in the building of foundation models for particle physics.
I Introduction
Foundation models are a new class of machine learning models which are trained on broad datasets and problems and are able to generalize to a variety of downstream tasks and datasetsĀ Bommasani etĀ al. (2022). Large-language models (LLMs) such as BERTĀ Devlin etĀ al. (2019), BARTĀ Lewis etĀ al. (2019), GPT-3Ā Brown etĀ al. (2020), and LLaMAĀ Touvron etĀ al. (2023) are examples of foundation models for text data while DALL-EĀ 2Ā Ramesh etĀ al. (2022) is an example of an image-based model.
The benefits of a foundation model for particle physics data would be significant: While machine learning models developed so far typically outperform classical approaches (and often by a large margin)Ā Kasieczka etĀ al. (2019a); Karagiorgi etĀ al. (2022), available statistics for training these models is a constant issueĀ Macaluso and Shih (2018); Qu etĀ al. (2022a). It becomes even more extreme in the case of searches for rare processes, where only a small fraction of simulated events might pass pre-selection criteria. Foundation models on the other hand are pre-trained and need fewer examples to be fine-tuned for a specific taskĀ Vigl etĀ al. (2024).
Beyond improving the physics performance by e.g. increasing selection accuracy of classification tasks, foundation models also address the other major issue currently facing particle physics: limited computing resources in the face of an ever increasing amount of dataĀ Albrecht etĀ al. (2019); Boehnlein etĀ al. (2022). This problem has already spawned the development of e.g. increasingly high fidelity models for the simulation of calorimeters, as well as techniques for the speeding up of other Monte Carlo simulationsĀ Paganini etĀ al. (2018); Buhmann etĀ al. (2021, 2023a); Adelmann etĀ al. (2022); Badger etĀ al. (2023); Hashemi and Krause (2023); Butter etĀ al. (2019); deĀ Oliveira etĀ al. (2017). By allowing the re-use of models across datasets and tasks, foundation models will play an important role in reducing this computational burden. Note that this effect will be further compounded by the potential for optimization in computing operations from one model used across multiple tasks.
Finally, using closely related architectures across different tasks inside one experiment, across experimental collaborations, and in the exchange with the theory community will make results easier to re-interpretĀ ara (2024); Bieringer etĀ al. (2024).
These potential benefits have inspired research into proto-foundation models suitable for particle physics. For example, Dillon etĀ al. (2022a); Favaro etĀ al. (2023); Dillon etĀ al. (2023); Park etĀ al. (2023); Dillon etĀ al. (2022b) investigated how known physical symmetries could be used to learn powerful embeddings of jet data, Benato etĀ al. (2022) showed the versatility of graph-based message passing networks for datasets from different domains of physics, Liu etĀ al. (2023); Salamani etĀ al. (2023) demonstrated conditioning generative models on the geometry of the detector to allow the simultaneous simulation of multiple detectors with one architecture, Dolan and Ore (2022); Beauchesne etĀ al. (2024) used meta-learning for mass-decorrelation and weak-supervision, and Qu etĀ al. (2022a) achieved state-of-the-art performance on the top tagging landscape dataset Kasieczka etĀ al. (2019b) by pre-training on a different datasetĀ Qu etĀ al. (2022b) and transferring the results.
Due to their flexibility demonstrated across language and other domains, transformers Vaswani etĀ al. (2017) are currently the most suitable candidate architecture for building foundation models for applications in particle physics. Taking inspiration from LLMs where sentences are generated autoregressively, recent efforts have demonstrated success with autoregressive generation of particle physics data, for example using the encoder part of the transformer to generate and jets Finke etĀ al. (2023) and using the decoder part of the transformer to generate +jets events Butter etĀ al. (2023). Ref. Heinrich etĀ al. (2024) demonstrated how a transformer backbone can be pre-trained using a BERT-like scheme where the model is trained to predict masked out jet constituents, resulting in an improvement of the performance when fine-tuning the backbone (with a new classification head) for jet tagging, especially at small training dataset sizes. Furthermore, Huang etĀ al. (2024) showed how a tokenized detector representation can be used in a BERT-like model for track reconstruction.
In this work, we will explore whether an autoregressive Generative Pretrained Transformer (GPT) model can be used as a foundation model for jet physics. However, the standard GPT constructions are not built to deal with continuous input data, but rather tokenized data. As point clouds are the most versatile representation of physics data Komiske etĀ al. (2019); Buhmann etĀ al. (2023b); Kasieczka etĀ al. (2019a); Buhmann etĀ al. (2023c, a) and can incorporate both event level information, jet substructure, and even low-level detector signals, finding a suitable input transformation for point clouds to tokens is the most pressing problem. Various tokenization strategies have been explored, for example using a simple mapping based on binning the input space in Finke etĀ al. (2023), a Gaussian mixture model in Butter etĀ al. (2023), and using an additional conditional embedding network in Heinrich etĀ al. (2024).
Here, we follow the conditional tokenization strategy from vanĀ den Oord etĀ al. (2018); Bao etĀ al. (2022); Heinrich etĀ al. (2024), but first take a step back to verify the quality and trade-offs involved in building these tokens. This will allow us to formulate quality measures to choose a suitable tokenization model, leading to an increase in codebook size from 512 tokens inĀ Heinrich etĀ al. (2024) to 8192 tokens.
Using this representation, we will first demonstrate training a generative model for jets as tokens in an unsupervised way for the JetClass Qu etĀ al. (2022b) dataset. Compared to Finke etĀ al. (2023), the core of our architecture is a transformer-decoder, not a transformer-encoder.
Finally, this allows us to test whether the information encoded in a model that was trained to generate jets can also be transferred to the task of classifying them. Observing such a transfer ability across different classes of tasks ā as opposed to transfer between different classification or generation problems ā would be a crucial ingredient to building foundation models for physics data, and has not yet been achieved. A graphical representation of this approach is provided in FigureĀ 1. As this is the first prototype of a model to tackle all tasks with jets in particle physics, it is named OmniJet-.
The rest of the paper is organized as follows: SectionĀ II introduces the data as well as the tokenization approach, the generative architecture, and the transfer learning strategy. Next, SectionĀ III shows the results of the tokenization study, the generative performance, as well as tests of the transfer learning capabilities of the model. Finally, SectionĀ IV summarizes the results and provides a brief outlook.
II Methods and Dataset
II.1 Dataset
All studies are performed using the JetClass datasetĀ Qu etĀ al. (2022b), originally introduced inĀ Qu etĀ al. (2022a). It contains both jet-level and constituent-level features for ten different types of jets initiated by gluons and quarks (), top quarks (, subdivided by their decay mode into and ) , as well as , , and (, , , , and ) bosons.
Events are simulated using MadGraph5_aMC@NLO Alwall et al. (2014) with parton showering and hadronization done by Pythia Sjöstrand et al. (2015). A simplified detector simulation implemented in Delphes de Favereau et al. (2014) using the CMS detector The CMS Collaboration (2008) card is performed. Constituents are clustered into jets using the anti- algorithm Cacciari et al. (2008) with a distance parameter of .
Jets are selected if they have a transverse momentum of and a pseudorapidity of . Additionally, truth-level matching is performed for all classes except and only jets that contain all the decay products of the boson or top quark are included. The resulting dataset contains 100M jets for training, 5M jets for validation, and 20M jets for testing.
In this work, only the kinematic information per particle (, , ) is used while the particle mass is approximated as zero. Next, the azimuth angle and the pseudorapidity are pre-processed to be relative to the jet axis111The difference in is signed and rectified to through . We handle those calculations using the scikit-hep/vectorĀ Schreiner etĀ al. (2023) and scikit-hep/awkwardĀ Pivarski etĀ al. (2024) libraries. :
Finally, we apply the cuts and to remove a very small fraction of low-energy constituents at the periphery and use up to 128 particles per jet.
II.2 Jet constituent token creation
We explore three kinds of tokenization approaches: binned, conditional, and unconditional tokenization. In the binned approach Finke etĀ al. (2023), the space of input features is subdivided using a regular grid in all dimensions (e.g. a 21x21x21 grid in three dimensions) and the cells in this grid are enumerated, resulting in one token per cell.
In the unconditional approach, each constituent is tokenized individually using a non-linear mapping, whereas in the conditional approach constituents are encoded and decoded conditioned on each other. We use a Vector Quantized Variational AutoEncoder (VQ-VAE)Ā vanĀ den Oord etĀ al. (2018); Bao etĀ al. (2022); Heinrich etĀ al. (2024); Huh etĀ al. (2023) to create a discrete set of jet constituent tokens both for conditional and unconditional tokenization.
The input features for the VQ-VAE are the , and values of the jet constituents. For the conditional tokenization, we use a transformer for both the encoder and the decoder of the VQ-VAE, whereas a simple multi-layer perceptron (MLP) is used for the unconditional tokenization. Details about the different VQ-VAE models used in our studies, as well as details about the preprocessing of the input features can be found in subsectionĀ A.1.
II.3 Transformer backbone
The core of OmniJet- is a transformer backbone based on the GPT transformer decoder model first introduced inĀ Radford etĀ al. (2018). However, since jet constituents are permutation invariant, we do not employ the positional encoding usually used in LLMs. As input, the transformer backbone receives the generated tokens from the VQ-VAE, complemented with a start and stop token. A jet with constituents is then represented as
where are the tokens.
The transformer backbone itself consists of an embedding layer followed by a series of GPT blocks. Each GPT block contains a multihead attention block, followed by a residual addition, layer normĀ Ba etĀ al. (2016), two linear layers with a ReLU in between, another residual addition and a final layer norm. Since this is an autoregressive model, a causal mask is passed together with the input data to the multihead attention block to prevent the model from seeing future tokens. The architecture is illustrated in FigureĀ 2.
The output from the transformer backbone is passed to a task specific head, either for generation or classification. The generative head is a single linear layer, while the classification head consists of a linear layer followed by ReLU, a sum over the constituent dimension, and a last linear layer with softmax activation function. The model is trained with heads in the multi-head attention block and GPT blocks. No dropout is used.
Once the generative model, i.e. the transformer backbone together with the generative head, has been trained on the tokenized data, it can generate new data autoregressively. The model has learned the probability distribution for a token , given a sequence of tokens:
The model is provided with the start token, and then samples this distribution to sequentially generate new tokens. Generation is repeated until either the stop token is generated or the maximum sequence length (which is set to be equal to 128) is reached. The generated token sequences are then fed to the VQ-VAE decoder, which maps them into physical space for further evaluation.
The classification task can be performed either from scratch, using randomly initialized weights for both the transformer backbone and the classification head, or by fine-tuning the generative model. In the fine-tuning case, the initial weights of the transformer backbone are loaded for from the generative model.
III Results
III.1 Token quality
We first inspect how well the tokens cover the space. An illustration of the conditional and unconditional tokens in physical space (i.e. their corresponding , and values) is shown in FigureĀ 3 for the different tokenization approaches and different codebook sizes. In the unconditional case, as well as in the binning approach, the reconstruction of a token is always the same, independent of the other tokens in the jet, leading to discrete points in physical space. In the conditional case, however, the reconstruction of a token is by construction affected by the other tokens inside this jet. To visualize the spread of each conditional token in physical space, we reconstruct each token 500 times conditioned on 50 randomly chosen tokens. Each of those reconstructions is shown in the scatter plots in FigureĀ 3, where the different reconstructions of the same token are drawn in the same color. We notice that the reconstruction of each token only changes slightly when conditioned on other tokens. Thus, the 500 different reconstructions of a conditional token show up in FigureĀ 3 as a blob in physical space. This already shows that the conditional tokenization allows to cover a significantly larger volume in reconstruction space, while the unconditional tokens can only be reconstructed to distinct points in reconstruction space. We note that our approach of reconstructing each token 500 times conditioned of randomly chosen other tokens not necessarily represents the reconstructed values of actual jet constituents, as it is possible that those combinations of tokens would not appear for real jets. However, this illustrates the overall behavior of how much the reconstruction of a token can change due to the conditioning on the other tokens.
Next, we consider distributions at the jet level to judge the quality of the tokenization. For this and the following studies, jets are mapped into token space, and then mapped back to physical space to quantify the loss in information. FigureĀ 4Ā (left) shows the jet mass combined for all classes in the dataset, as was done inĀ Heinrich etĀ al. (2024). As observed there, the conditional tokenization with 512 tokens already shows a good agreement between initial and reconstructed mass, with worse performance for the unconditional tokenization.
However, this inclusive distribution might hide differences at the level of individual classes and jets. We therefore also consider the difference in mass for jets before and after tokenization and reconstruction FigureĀ 4Ā (center) for jets. The unconditional tokenization leads to a systematic shift of approximately 15Ā GeV, while the conditional tokenization is well centered at zero. Increasing the codebook size from 512 to 8192 tokens substantially improves the resolution. This behavior is even more pronounced when considering the -subjettinessĀ Thaler and VanĀ Tilburg (2011) ratio in FigureĀ 4Ā (right). Both conditional and unconditional tokenization with 512 tokens results in shifted distributions, while the larger codebook size of 8192 recovers a peak close to zero. Furthermore, while the mass resolution of the conditional tokens is already centered close to zero when using a codebook size of 512, the width of the distribution improves drastically from Ā GeV to Ā GeV when moving from a codebook size of 512 to a codebook size of 8192, where corresponds to the standard deviation obtained from fitting a normal distribution to the mass resolution histograms. A similar behavior can be observed for other classes, where in some cases, depending on the jet observable and the jet type, the effect is even more extreme. Finally, the binning approach with a 21x21x21 linear bins in the input features of our VQ-VAE comes close to the mass resolution of the conditional tokens with a codebook size of 8192, while the resolution of the subjettiness ratio is notably worse. Moreover, while this binning approach with 9162 tokens leads to reasonable resolution of the jets shown in FigureĀ 4, we found quite drastic mismodeling for and jets with such small codebook sizes222As expected, the resolution of the binning approach automatically leads to good resolution when the codebook size (i.e. the the number of bins) is increased to a sufficiently large number. We found that around tokens (corresponding to a 40x40x40 grid) offer similar resolution as conditional tokenization with a codebook size of 8192. . The distributions and the corresponding resolutions of the jet mass, jet , as well as the subjettiness ratios and are shown for all ten jet types individually in AppendixĀ B. Overall, the highest fidelity is achieved by conditional tokenization with a marked improvement from increasing the codebook size from 512 to 8192.
Finally, we quantify the information loss that comes with the tokenization by training multi-class classifiers to distinguish between the ten jet types present in the dataset. The classifiers are trained once with the original inputs, and once with the inputs after undergoing tokenization and subsequent reconstruction back into physical space. This procedure allows a direct comparison of how the loss in resolution affects reconstruction performance. We utilize two standard classifier architectures: DeepSetsĀ Zaheer etĀ al. (2018); Komiske etĀ al. (2019) (i.e. without particle interactions) and TransformerĀ Vaswani etĀ al. (2017); Shleifer etĀ al. (2021) (i.e. with particle interactions) and perform this study for four different codebook sizes from 512 to 8192 tokens for the conditional tokenization approach. Details about the classifier trainings can be found in subsectionĀ A.2. Note that this approach is similar in spirit to the classifier metric proposed inĀ Krause and Shih (2023); Das etĀ al. (2023) but tests a multi-class classifier trained on these samples individually, as opposed to judging how well a classifier might distinguish original and reconstructed samples. This is necessary as e.g. points at fixed positions would be distinguishable from the original with close to perfect accuracy, rendering the test less useful.
The resulting classifier accuracy for the two different architectures is shown in FigureĀ 5. As seen in previous studies of resolution, we observe an increase of token quality as we increase the size of the VQ-VAE codebook. Furthermore, we see that the classifier performance starts to plateau with codebook sizes larger than 4096. However, even at the largest codebook size, a gap to the performance on original particles remains, motivating future work into building more accurate tokenization schemes.
For the remaining studies we will utilize a codebook size of 8192 with conditional tokens as this leads to the overall best performance and fidelity.
III.2 Jet generation
After training the transformer backbone with the generative head, it can be used for autoregressive generation as described in subsectionĀ II.3. The model was trained on three separate datasets: only, only, and and combined. This section will describe the combined version, since this is the model that is used for transfer learning. For a discussion of single-jet type generative results, including a comparison to the EPiC-FM method ofĀ Birk etĀ al. (2023), see appendix AppendixĀ C. 48Ā 000 events were generated from the combined model. These events contain tokens, which are then decoded back to physical space using the VQ-VAE decoder.
A comparison to reconstructed JetClass tokens can be seen in FigureĀ 6. We observe that in general the model is able to match the truth level tokens well. We note that the tail of the spectrum of both the generated constituents and the reconstructed JetClass tokens contains bumps distributed around discrete values, which is consistent with our inspection of the reconstruction space shown in FigureĀ 3.
In order to quantify the performance, a classifier test (see subsectionĀ A.5 for details) is performed to distinguish generated events from reconstructed JetClass tokens. The test results in an AUC score of 0.54.
III.3 Transfer learning from generation to classification
To evaluate the ability of the model to generalize from generating jets to classifying them, we focus on the task of hadronic top quark taggingĀ Kasieczka etĀ al. (2019a), i.e. distinguishing and jets on the JetClassĀ Qu etĀ al. (2022b) dataset. For this test, the Next-token prediction head is replaced by a Classification head while the transformer backbone is retained. We compare three training strategies: training the full architecture with randomly initialized weights (termed from scratch) which does not use transfer-learning and corresponds to the baseline, and two versions of fine-tuning the model obtained in the generative training step. In the Fine-tuning run, both the pre-trained backbone weights and the randomly initialized classification task head weights are allowed to float in the training, while in Fine-tuning (backbone fixed) only the classification task head is allowed to change.
The results of these training runs are presented in FigureĀ 7 as a function of the number of training examples provided to the model. We observe a significant gain in classification accuracy of both fine-tuning approaches compared to the baseline, leading to up to 15 percentage-points higher accuracy for small number of training jets, and outperforming by a few percentage-points at the highest training sample size. The difference between the two fine-tuning strategies is relatively small, with the more open training performing slightly better. Put differently, the generative pre-trained model achieves the same accuracy around 84% with 100 training examples for which the model that is trained from scratch requires 10ā000 examples.
IV Conclusion
Foundation models for physics data are an enticing promise: Trained on large amounts of data and tasks, they are expected to easily generalize to any down-stream problem, saving countless hours of human and compute time. In this paper we have taken crucial steps towards the creation of such models.
First, we expect learned representations of data to play a key role as inputs to foundation models. Representations might be continuous and rely on symmetries or learn a mapping to a discrete space as done here with tokenization. Note that while using data raw ā i.e. without prior mapping into a representation space ā might be possible when only considering a narrow range of similar datasets, it is inherently limiting when data from different sources or with different initial dimensionalities are to be considered.
Whatever representation is used, it will be important to understand and minimize the loss of information inherent in this transformation. This problem is especially important for downstream uses such as classification and regression tasks, as the loss of information can directly limit the achievable accuracy or resolution. This work introduced a set of criteria ā both distribution and classifier based ā that can be used to assess the quality of any representation.
Using these metrics, we found a marked increase in the resolution of relevant observables like mass and jet substructure by using a codebook size of 8192 with conditional tokenization over binning-based approaches, unconditional tokenization, and conditional tokenization with smaller codebooks. An additional classifier test further confirmed this observation.
Next, we demonstrated the training of an autoregressive generative model for jet constituents, specifically for and jets from the JetClass datasetĀ Qu etĀ al. (2022b). The generated distributions agree well with the ground truth, both for global jet kinematics, jet substructure, and individual constituent features. We note that while our model is the first token-based generator of JetClass-like examples, more extensive studies of its generative fidelity when increasing the feature-space and detailed comparison to prior non-token-based results on this datasetĀ Birk etĀ al. (2023) are left for future work.
Most importantly, we report the generalization capability of OmniJet- from learning to generate jets in an unsupervised way, to the supervised classification between and jets. Overall, the fine-tuned model outperformed training from scratch for all values of training examples, often by a significant margin. For example, for 1000 training jets, the fine-tuned model achieves an accuracy of approximately 90%, compared to around 74% for the freshly initialized model. While the two approaches seem to converge, even at the highest training size of 2 million jets the fine-tuned approach maintains a lead of a few percentage points. Finally, it even provides a non-trivial classification accuracy of 84%, even when trained only on 100 jets, emphasizing the value of foundation models for problems with few available labeled training examples. While other types of transfer have been demonstrated previously, this is the first time that the unification across the two big classes of tasks ā classification and generation ā has been achieved.
Of course, this work is only one step in building overarching foundation models. While it is the first model that achieves both classification and generation, it is not the most performant for either of these tasks. However, strategies to increase the performance exist and will be integrated. For example, the representation quality needs to be improved, possible gains from masked pre-trained have to be evaluated, architecture and training data need to be scaled up, and more extensive studies of the generalization capabilities, including training and testing on additional tasks, performed. In the medium term, strategies need to be found to align diverse datasets as well as to integrate pre-trained foundation models in community workflows. Nevertheless, the potential benefits in physics performance and compute efficiency glimpsed at in this and other works makes this a worthy endeavor.
Acknowledgements
We thank David Shih, Michael KrƤmer, Michael Kagan, Frank Gaede, Sarah Heim, and Judith Katzy for stimulating discussions of foundation models for physics data and Erik Buhmann for valuable comments on the manuscript.
The authors acknowledge support by the Deutsche Forschungsgemeinschaft under Germanyās Excellence Strategy ā EXC 2121 Quantum Universe ā 390833306, and under PUNCH4NFDI ā project number 460248186. This research was supported in part through the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany.
Appendix A Model details and hyperparameters
A.1 VQ-VAE for token creation
Both the and the values are scaled down by a factor ofĀ 3. The transverse momentum of the jet constituents is first transformed using the natural logarithm and subsequently shifted by -1.8. The tokenization was also done without the log transform of the , and was found to perform similarly. However, the logarithm transformation has the advantage that it automatically avoids negative values, which is why we choose to use the log-transformed . The conditional and unconditional tokenization only differ in the architecture of the encoder and decoder of the VQ-VAE.
Training for the VQ-VAE is implemented in pytorchĀ Paszke etĀ al. (2019) and pytorch-lightningĀ Falcon and team (2024).
The model architecture of the VQ-VAE encoder and decoder in the conditional tokenization approach is similar toĀ Heinrich etĀ al. (2024) with a different set of hyperparameters. We use 4 NormFormerĀ Shleifer etĀ al. (2021) blocks with an embedding dimension of 128 and 8 heads in the MHA for both the encoder and the decoder. We use the vqtorch libraryĀ Huh (2022); Huh etĀ al. (2023) to implement the vector quantization layer with the dimension of the codebook vectors set to 4.
The mean squared error (MSE) between the tensor of the initial particle features and the reconstructed features is used as the task loss . The total loss is then set to
with . An affine transformation is used for a joint transformation of all codes and dead codes are replaced with a frequency of 10 iterations. The parameter which trades off the importance of updating the embeddings from the encoder and the code vectors is set to . Lastly, we use a synchronized update ruleĀ Huh etĀ al. (2023); Huh (2022) with .
In the unconditional approach, we use the same hyperparameters as outlined above, with the only difference that the architecture of the encoder and decoder is a simple MLP with 3 hidden layers of dimension 128 and ReLU activation function.
All VQ-VAE models are trained on all 10 classes of the JetClass datasetĀ Qu etĀ al. (2022b).
A.2 Classifiers for token quality evaluation
The DeepSetsĀ Zaheer etĀ al. (2018); Komiske etĀ al. (2019) classifier consists of a per-particle MLP with shared weights across all particles inside the jet with 3 hidden layers of dimension 100, 100 and 256. The output of the network is then aggregated with a sum and fed into another MLP with 3 hidden layers of dimension 100 followed by a 10-dimension output layer with softmax activation function.
The Transformer classifier consists of 5 NormFormerĀ Shleifer etĀ al. (2021) blocks, followed by two class-attention blocks with a class token as query, inspired by the ParTĀ Qu etĀ al. (2022a) architecture. The output of the last class-attention block is fed into a MLP with two hidden layers of dimension 128, followed by a softmaxed 10-dimensional output layer.
The classifiers are trained with the AdamWĀ Loshchilov and Hutter (2019) optimizer with a maximum learning rate of 0.005 (0.001) for the DeepSets (Transformer) classifier and weight decay 0.01. The learning rate first linearly increased from 0.002 (0.0005) during the first 4 training epochs, after which it is linearly decreased to the initial learning rate over 20 epochs and then linearly decreased to a final learning rate of 0.001 (0.0003), following the one-cycle learning rate schedule Smith (2018).
The classifiers for those token quality evaluations are trained on 10M jets from the JetClass datasetĀ Qu etĀ al. (2022b).
A.3 Transformer backbone
When training the transformer backbone, cross entropy is used as a loss function and AdamĀ Kingma and Ba (2017) with a learning rate of 0.001 as optimizer. The model had access to 10M jet events and 10M jet events. Note that this means that the model trained on these two jet types combined had access to twice as much data. All versions were trained for 30 epochs, and the model state with the lowest validation loss was chosen for the further analysis.
A.4 Transfer learning
To perform the transfer learning from the generative task to the classification task, we change the head of the OmniJet- model to the classification head and load the weights of the backbone trained for the generative task. We explore two variations of fine-tuning the pre-trained backbone to the classification task: training all weights of the model with the same learning rate (referred to as Fine-tuning in subsectionĀ III.3) and keeping the weights of the backbone fixed at the state obtained from the generative model (referred to as Fine-tuning (backbone fixed) in subsectionĀ III.3). For the From scratch trainings we start the training with randomly initialized weights of the whole model. The training is performed with the AdamWĀ Loshchilov and Hutter (2019) optimizer with a constant learning rate of 0.0001 and weight decay 0.01. Since those trainings, depending on the size of the training dataset, tend to show overfitting quite quickly, we stop those trainings when the validation loss does not improve for multiple epochs. The threshold of this early stopping is adjusted to the training dataset size, with a patience of 20 epochs for a training dataset of , and jets, a patience of 10 epochs for a training dataset size of and , and a patience of 5 for trainings with training jets. For each training dataset size we run 5 trainings with different random seeds and the epoch with the smallest validation loss is chosen for evaluation.
A.5 Classifier tests
In order to quantify the performance of the generative model, a classifier test using the structure of the DeepSets classifier from subsectionĀ A.2 is performed. In this case however, the 3 hidden layers of the particle MLP all have dimension 10.
A number of 48Ā 000 generated events are combined with equally many reconstructed tokens from the test set of JetClass. The two datasets are combined and shuffled, and a train/val/test split of 0.6/0.2/0.2 is used. The model is trained for 100 epochs with binary cross entropy loss and AdamĀ Kingma and Ba (2017) with learning rate 0.001 as optimizer. The model state with the lowest validation loss is chosen for evaluation. The resulting AUC scores are 0.54 for the model trained on and jets combined and 0.57 for the ones trained on single-type jets.
Appendix B Token quality
Additional plots of the jet mass, the jet transverse momentum, as well as the subjettiness ratios are shown in FigureĀ 8, FigureĀ 9, FigureĀ 11 and FigureĀ 10.
Appendix C Generative model trained on single-jet data
To test the generative performance, the generative model was also trained on single-jet type data ā 10M jets each of and ā separately. For these training, no tests of the task-transfer to classification were perfomed.
The result of the jet training is shown in FigureĀ 12, of the jet training in FigureĀ 13. In the case, we see a good agreement between the reconstructed tokens and the generated events. However, it seems to be somewhat more difficult for the model to accurately model for jets, which is also mirrored for this quantity in the combined model (see FigureĀ 6).
It is interesting to compare the result of OmniJet- with that of a different generative model. EPiC-FMĀ Birk etĀ al. (2023) was the first generative model trained on the JetClass dataset, utilizing flow matching and operating without tokenization. The result of the comparison can be seen in FigureĀ 14. The plots show the JetClass data, the reconstructed JetClass token from this work, the EPiC-FM generated events, and the OmniJet- generated events. We use the more challenging class for comparison.
The ratio plots under the main plots show the generated events compared to their respective truths: direct JetClass for EPiC-FM and Reconstructed JetClass tokens for OmniJet-. Hence, the ratios show how well the respective generative models learn to replicate their training data. In general, we see that both models are doing well. OmniJet- has a somewhat higher discrepancy in the tails of all distributions except for constituent and the number of constituents. The most difficult distribution is the constituent , with bumps in the tail, which could also be seen in FigureĀ 13.